2 moles of an ideal gas allowed to expand reversibly and isothermally at 300k frm pressure 1 atm to a prssur 0.1 atm.the change in gibbs energy?

Grade:12th pass

1 Answers

Abhishek Singh
7 years ago
i think you well know the equation dG = VdP-SdT
since the gas is expanding here in const temp. he SdT term is zero
so dG = VdP
you can intregrate dG to get gibbs energy change, by putting v = nRT/p
deltaG  = nRTln(P2/p1)
so answer is 2*8.314*300*ln(1/10) = -2*8.314*300*ln(10)